OpenStudy (anonymous):

simplify rational expression -15/3x-9 steps please!

OpenStudy (anonymous):

is it \[-\frac{15}{3x-9}\]?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

\[-\frac{15}{3x-9}=\frac{5}{3-x}\]